OpenStudy (unofficialllyy):

Help with 5 questions please?? :((

OpenStudy (unofficialllyy):

1. Given the triangle below, which of the following is a correct statement? A. cot angle B=6.5 B. csc angle C=3/4 C. cot angle C=8/5 D. csc angle B=4/3

OpenStudy (unofficialllyy):

@peachpi could you help?

OpenStudy (amtran_bus):

.